Hi everybody!
I just wanted to know, if you have a parallelogram or quadrilateral..or What ever,do you use 180 for the equal angles? Because heres what I did and I felt like i got the correct answer: PARALLELOGRAM:angle#1:138 angle#2:x angle#3:x angle#4:(3x+20)
First i set it up in an equation: 138=x+x+(3x+20)=180. Then i subtracted 138 from 138 and 180: 138-138+x+x+3x+20=180-138 and then I got 42. I combined like terms..i put it all together:5x+12=42. I subtracted 42-12 and got 30. Then i did 30/(divided)by 5x and got 6 as my final answer. :138+6+6=3x6+20=180..Is tht correct?..If not please explain why and how to do it correctly.^_^ please and Thankyou!!!

Answer by htmentor(1343)

First of all, the sum of all the angles of a quadrilateral is 360 degrees. Consider the special case of a square or rectangle.
In this case, we know all 4 angles are right angles, so the sum of all the angles = 90+90+90+90=360.
Secondly, if the angles given in the problem are 138,x,x,3x+20, then the figure cannot be a parallelogram.
In a parallelogram opposite sides are parallel, and opposite angles are equal.
If we solve for x, we get:
138 + x + x + 3x + 20 = 360
5x = 360 - 158
x = 40.4
This would give the following angles:
138,40.4,40.4,141.2
These angles cannot form a quadrilateral. If two of the angles are equal, then the other two must also be equal.